# Topological Data Analysis for Neural Network Analysis: A Comprehensive Survey

### 10 Dec 2023 v1

@@ 3.3.1 Mapper
Two seminal papers by Carlsson and Brüel Gabrielsson, published in 2019 and a follow-up in 2020, stand as early explorations of the use of Mapper to analyze internal representations of FCFNNs, specifically CNNs. These papers focus on the topology and distribution of convolutional filter weights within the same layer of CNNs.
@@
One of the visualization techniques proposed by Rathore et al. was the computation of activation images for the activations collected in each node and for its average. Activation images, given an activation vector or an average of them, are idealized images that would have produced the activations via an iterative optimization process proposed by Olah et al. (2017, 2018). Other statistical data proposed by Purvine et al. (2023) to describe Mapper nodes were: 1. Pie charts showing the composition of class labels in that node, associating to each activation vector the label of the image that generated the activation vector; 2. Node-wise purity, defined as $\alpha_{i}=c_{i}^{-1}$ where $c$ is the number of class labels in the Mapper node $i$; 3. Class-wise purity, defined as $\gamma_{y}=\left|J_{y}\right|^{-1}\sum_{x\in J_{y}}\beta_{x}$, where $J_{y}$ is the set of activations in the Mapper graph associated to the class $y$ and $\beta_{x}=\left|I_{x}\right|^{-1}\sum_{i\in I_{x}}\alpha_{i}$, with $I_{x}$ being the set of nodes in the Mapper graph containing the activation $x$.
@@ 3.3.2 Homology and persistent homology
While in previous works the topology of the activations was studied globally, Naitzat et al. (2020) studied the evolution of the activations layer by layer in FCFNNs for binary classification problems in an extensive set of experiments. For data samples $\mathcal{D}_{i}$ from one of the classes $i\in[2]$ at a time, the evolution of the (persistent) homology of $\mathcal{D}_{i}$ through the different layers is analyzed. In particular, for each example $x\in\mathcal{D}_{i}$ and each layer $l$ in an FCFNN $\mathcal{N}$, an activation vector $a_{x}^{l}$ is computed by taking the vector composed of the different activation values of the neurons given the example $x$ in the layer $l$, that is,
@@
The previous observations were contradicted by Wheeler et al. (2021). In this work, evolution of the topology of the activations through the different layers of FCFNNs for synthetic and real data was studied again. However, this time, the activations and distances between them were preprocessed and slightly modified depending on the experiments performed to build a point cloud. From such preprocessed point clouds, Vietoris–Rips persistence diagrams were computed using the usual Euclidean distance between vectors. Finally, instead of studying Betti numbers or persistence diagrams directly, Wheeler et al. computed persistence landscapes $\lambda^{l}_{k}=\left(\lambda^{l}_{k,1},\lambda^{l}_{k,2},\ldots\right)$ derived from the persistence diagrams for each layer for different homological degrees $k$.
@@
$H_{1,l,c,y}=\begin{cases}H^{\prime}_{1,l,c,y}&\text{if }\varepsilon_{1,l,c,y}\geq p,\\[2.84544pt] (1-\varepsilon_{1,l,c,y})\text{log}\left|\Omega_{y}\right|&\text{otherwise},\end{cases}$
@@
Let $\mathcal{N}$ be a neural network, and $v_{l}^{i},v_{l+1}^{j}\in V(G(\mathcal{N}))$ be two connected neurons in the directed neural network graph $G(\mathcal{N})$. Formally, the relevance of the directed edge $(v_{l}^{i},v_{l+1}^{j})$ connecting $v_{l}^{i}$ and $v_{l+1}^{j}$ in the output calculation is defined to be the linearly rectified weight of the edge normalized by the sum of the linearly rectified weights of the edges pointing to $v_{l+1}^{j}$, that is,
@@
$S_{v}=\sum_{((v,w)\in E(G(\mathcal{N})))}\!\!\!\!\!\!\!\pi_{w,v}\cdot\delta_{w},\quad\pi_{w,v}=\frac{\left|W_{w,v}\right|}{\sum_{(u,w)\in E(G(\mathcal{N})))}\left|W_{w,u}\right|},\quad\delta_{w}=\begin{cases}1&\text{ if }w\in V_{L},\\ S_{w}&\text{ otherwise, }\end{cases}$
@@
Layerwise topological complexities have also been studied as a function of the weights. Rieck et al. (2019) proposed to study the so-called neural persistence. Given a non-output layer $l$ of a neural network $\mathcal{N}$, its neural persistence is
@@
To compare different layers from the network, Rieck et al. proposed to normalize the neural persistence by dividing it by
@@
that is, an upper bound of the neural persistence, obtaining the normalized neural persistence $\tilde{\text{NP}}_{l}(\mathcal{N})$ for each layer $l$. Averaging normalized neural persistences over all the layers of the network, a global topological complexity measure $\bar{\text{NP}}(\mathcal{N})$ is obtained.
For simple FCFNN networks trained on MNIST, neural persistence was able to distinguish clearly between properly and badly trained networks, for which the values of the neural persistence of properly trained ones were consistently higher than for their badly trained counterparts. Furthermore, it was observed that different regularization techniques augmented the mean neural persistence with respect to the values of regular trained networks, suggesting that for a fixed architecture, the higher the neural persistence, the better the generalization capabilities of the network. This was exploited as an early stopping criterion for training neural networks, where the training process is completed when the mean neural persistence $\bar{\text{NP}}(\mathcal{N})$ stops increasing significantly. This early stopping criterion was found to be competitive with other early stopping criteria but without the need for a validation dataset, whose use is not always possible due to data scarcity.
Girrbach et al. deepened more into the properties of neural persistence. On the one hand, they discovered tighter bounds than those originally presented by Rieck et al. (2019) for the value of neural persistence. On the other hand, they also observed that there is a close relationship between the variance of the learned weights of deep learning models and the neural persistence, questioning the value of the latter with respect to this simpler measure to study the properties of the neural network, arguing that this variance may be similarly useful for the applications showcased by Rieck et al..
@@ 3.4 Training dynamics and loss functions
One of the fundamental problems in deep learning is, given a learning problem, an architecture $a$, a loss function $\mathcal{L}$, and a training algorithm $\mathcal{A}$, to determine if the training algorithm $\mathcal{A}$ is capable of finding a neural network $\mathcal{N}$ with architecture $a$ that minimizes the empirical risk ^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}$ for the learning problem. Usual training algorithms perform a gradient descent, following a path in the space of parameters of the neural network. This path is then heavily affected by the connectivity of the loss graph and by the presence of local valleys, in which the gradient descent algorithm can get stuck.
@@
A bad local valley is simply a local valley in which the target $\mathfrak{L}$ cannot be arbitrarily close to the lower bound of the convex function inducing the target function, which is also a lower bound of $\mathfrak{L}$ but does not depend on any concrete neural network. Bad local valleys are harmful to the optimization problem because if the training process enters them when optimizing the target function, the parameters obtained at the end of the learning process are provably not optimal. Nguyen proved that, for activation functions as before that also satisfy that there are no nonzero coefficients $(\lambda_{i},a_{i})_{i=1}^{p}$ with $a_{i}\neq a_{j}$ for all $i\neq j$ such that $\varphi(x)=\sum_{i=1}^{p}\lambda_{i}\varphi(x-a_{i})$ for all $x\in\mathbb{R}$, FCFNNs with a layer $l\in[L-1]$ satisfying $N_{l}>\left|\mathcal{D}_{\text{train}}\right|$ and $N_{i}>N_{i+1}$ for $i\in\{l+1,\ldots,L\}$ induce target functions $\mathfrak{L}$ with no bad local valleys and, if $l\leq L-2$, then every local valley of $\mathfrak{L}$ is unbounded. This implies that from any initial parameter, there is a continuous path on which the loss function is nonincreasing from it to a point that is arbitrarily close to the infimum of the loss.
@@
$\sup_{\theta\in\Theta}\left|\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\phi_{\mathcal{N}_{\theta}})-\mathcal{R}(\phi_{\mathcal{N}_{\theta}})\right|\leq 2B\sqrt{\frac{(\text{dim}_{\text{PH}}\Theta+1)\log^{2}(mL^{2})}{m}+\frac{\log(7M/\gamma)}{m}}$
@@ 4 Challenges, future directions, and conclusions
In this survey we have seen many examples of how topological data analysis, and particularly (persistent) homology and Mapper, can be applied to study the properties of neural networks. In Section 3.1, we saw how to compute exactly two different homology groups that yielded different information about a neural network only by taking its graph without weights. However, we noticed that the ranks of these homology groups were invariant to many important structural properties of the neural network, such as the order of the layers. Also, their values were simply a combination of the width and the depth of the graph, not capturing much information about the network. To obtain better insight into the properties of neural networks, more fine-grained homology theories for directed graphs are needed. A promising line of work would be to study the persistent path homology (Chowdhury and Mémoli, 2018) of graphs augmented with the weights associated to each edge, as in Section 3.3.
@@
In Section 3.2, we saw how topology was useful in recovering the topology of the decision regions and boundaries, with applications especially in model selection. Thanks to the differentiability theory of persistence diagrams, we saw how to improve these regions by simplifying them. In addition, we discussed how the topology was useful for capturing the quality, measured in different ways, of generative models. Particularly interesting is the study of the disentanglement of generative models: being able to decouple the different sources of variation could lead to a better control and quality of the networks’ outputs. We find that refining the article by Zhou et al. (2021) and using its measures to regularize generative neural networks could be an interesting approach to improving neural network design using TDA.
@@
In Section 3.3 we analyzed the methods studying the neural network parameters and activations. We split the section into the articles using Mapper and (persistent) homology, because of the differences in their approaches. In the first case, we found both studies of the parameters and of the activations. In the first case, we saw how Carlsson and Brüel Gabrielsson found very interesting connections between the topologies of the space of natural images and the space of weights of some layers in convolutional networks, among many other interesting topology configurations of the convolutional weights. Also, we saw how Gabella studied the Mapper graphs of the evolution of the weights during training, finding a very interesting pattern in the weight distribution of one of the architectures tried in his experiments: the weights seemed to be distributed in a surface! A closer look at the topology of the weight evolution for a higher variety of architectures must be taken to realize if the weights share common patterns in similar problems. This could potentially enable leveraging this structure, as proposed in Carlsson and Brüel Gabrielsson (2020), for instance. For activations, we found that the general approach proposed by TopoAct (Rathore et al., 2021) was incredibly useful for understanding the internal behavior of neural networks, with the many ramifications presented in the survey.
@@
These computational problems motivate the development of more suitable algorithms to compute persistence diagrams or alternative invariants that also capture information of neural networks. Although powerful tools for computing Vietoris–Rips persistence diagrams are available, such as Ripser (Bauer, 2021), Ripser++ (Zhang et al., 2020), Gudhi (The GUDHI Project, 2015), SLINK (Sibson, 1973), algorithms that parallelize computations are of special interest due to the increasing availability of powerful hardware and software methods to perform distributed computing. Also, approximating known filtrations by means of new filtrations, as in (Sheehy, 2013), or by means of machine learning models, as in Montufar et al. (2020).
In conclusion, we have seen that the use of TDA applied to studying neural networks is an exciting path that can be useful in many scenarios. TDA, although computationally expensive, is a tool that has been shown to be connected to many interesting properties of neural networks such as generalization in classification problems or entanglement and latent space quality in generative models. Moreover, we have seen that there are many ways in which TDA can be used in applied scenarios, converting TDA into a potential essential tool for deep learning practitioners.

### 03 Jan 2024 v2

@@ 3.3.1 Mapper
Two seminal papers by Carlsson and Brüel Gabrielsson  published in 2019 and a follow-up in 2020, stand as early explorations of the use of Mapper to analyze internal representations of FCFNNs, specifically CNNs. These papers focus on the topology and distribution of convolutional filter weights within the same layer of CNNs.
@@
One of the visualization techniques proposed by Rathore et al. was the computation of activation images for the activations collected in each node and for its average. Activation images, given an activation vector or an average of them, are idealized images that would have produced the activations via an iterative optimization process proposed by Olah et al. (2017, 2018). Other statistical data proposed by Purvine et al. (2023) to describe Mapper nodes were: 1. Pie charts showing the composition of class labels in that node, associating to each activation vector the label of the image that generated the activation vector; 2. Node-wise purity, defined as $\alpha_{i}=c_{i}^{-1}$ where $c_{i}$ is the number of class labels in the Mapper node $i$; 3. Class-wise purity, defined as $\gamma_{y}=\left|J_{y}\right|^{-1}\sum_{x\in J_{y}}\beta_{x}$, where $J_{y}$ is the set of activations in the Mapper graph associated to the class $y$ and $\beta_{x}=\left|I_{x}\right|^{-1}\sum_{i\in I_{x}}\alpha_{i}$, with $I_{x}$ being the set of nodes in the Mapper graph containing the activation $x$.
@@ 3.3.2 Homology and persistent homology
While in previous works the topology of the activations was studied globally, Naitzat et al. (2020) studied the evolution of the activations layer by layer in FCFNNs for binary classification problems in an extensive set of experiments. For data samples $\mathcal{D}_{i}$ from one of the classes $i\in[2]$ at a time, the evolution of the (persistent) homology of $\mathcal{D}_{i}$ through the different layers is analyzed. In particular, for each example $x\in\mathcal{D}_{i}$ and each layer $l$ in an FCFNN $\mathcal{N}$, an activation vector $a_{x}^{l}$ is computed by taking the vector composed of the different activation values of the neurons given the example $x$ in layer $l$, that is,
@@
The previous observations were contradicted by Wheeler et al. (2021). In this work, evolution of the topology of the activations through the different layers of FCFNNs for synthetic and real data was studied again. However, this time, the activations and distances between them were preprocessed and slightly modified depending on the experiments performed to build a point cloud. From such preprocessed point clouds, Vietoris–Rips persistence diagrams were computed using the usual Euclidean distance between vectors. Finally, instead of studying Betti numbers or persistence diagrams directly, Wheeler et al. computed persistence landscapes $\lambda^{l}_{k}=\big{(}\lambda^{l}_{k,1},\lambda^{l}_{k,2},\ldots\big{)}$ derived from the persistence diagrams for each layer for different homological degrees $k$.
@@
$H_{1,l,c,y}=\begin{cases}H^{\prime}_{1,l,c,y}&\text{if }\varepsilon_{1,l,c,y}\geq p,\\[2.84544pt] (1-\varepsilon_{1,l,c,y})\log\left|\Omega_{y}\right|&\text{otherwise},\end{cases}$
@@
Let $\mathcal{N}$ be a neural network, and $v_{l}^{i},v_{l+1}^{j}\in V(G(\mathcal{N}))$ be two connected neurons in the directed graph $G(\mathcal{N})$. Formally, the relevance of the directed edge $(v_{l}^{i},v_{l+1}^{j})$ connecting $v_{l}^{i}$ and $v_{l+1}^{j}$ in the output calculation is defined to be the linearly rectified weight of the edge normalized by the sum of the linearly rectified weights of edges pointing to $v_{l+1}^{j}$, that is,
@@
$S_{v}=\!\!\!\!\!\!\!\!\sum_{(v,w)\in E(G(\mathcal{N}))}\!\!\!\!\!\!\!\!\!\pi_{w,v}\cdot\delta_{w},\quad\pi_{w,v}=\frac{\left|W_{w,v}\right|}{\sum_{(u,w)\in E(G(\mathcal{N}))}\left|W_{w,u}\right|},\quad\delta_{w}=\begin{cases}1&\text{ if }w\in V_{L},\\ S_{w}&\text{ otherwise, }\end{cases}$
@@
Layerwise topological complexities have also been studied as a function of the weights. Rieck et al. (2019) proposed to study the so-called neural persistence. Given a non-output layer $l$ of a neural network $\mathcal{N}$, its neural persistence is defined as
@@
To compare different layers from the network, Rieck et al. proposed to normalize neural persistence by dividing it by an upper bound of the neural persistence,
@@
obtaining the normalized neural persistence $\widetilde{\text{NP}}_{l}(\mathcal{N})$ for each layer $l$. Averaging normalized neural persistences over all the layers of the network, a global topological complexity measure $\overline{\text{NP}}(\mathcal{N})$ is obtained.
For simple FCFNN networks trained on MNIST, neural persistence was able to distinguish clearly between properly and badly trained networks, for which the values of the neural persistence of properly trained ones were consistently higher than for their badly trained counterparts. Furthermore, it was observed that different regularization techniques augmented the mean neural persistence with respect to the values of regular trained networks, suggesting that for a fixed architecture, the higher the neural persistence, the better the generalization capabilities of the network. This was exploited as an early stopping criterion for training neural networks, where the training process is completed when the mean neural persistence $\overline{\text{NP}}(\mathcal{N})$ stops increasing significantly. This early stopping criterion was found to be competitive with other early stopping criteria but without the need for a validation dataset, whose use is not always possible due to data scarcity.
Girrbach et al. deepened more into the properties of neural persistence. On the one hand, they discovered tighter bounds than those originally presented by Rieck et al. (2019) for the value of neural persistence. On the other hand, they also observed that there is a close relationship between the variance of the learned weights of deep learning models and the neural persistence, questioning the value of the latter with respect to this simpler measure to study the properties of the neural network, arguing that this variance may be similarly useful for the applications showcased by Rieck et al.
@@ 3.4 Training dynamics and loss functions
One of the fundamental problems in deep learning is, given a learning problem, an architecture $a$, a loss function $\mathcal{L}$, and a training algorithm $\mathcal{A}$, to determine if the training algorithm $\mathcal{A}$ is capable of finding a neural network $\mathcal{N}$ with architecture $a$ that minimizes the empirical risk ^$\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}$ for the learning problem. Usual training algorithms perform a gradient descent, following a path in the space of parameters of the neural network. This path is then heavily affected by the connectivity of the loss graph and by the presence of local valleys, in which the gradient descent algorithm can get stuck.
@@
A bad local valley is a local valley in which the target $\mathfrak{L}$ cannot be arbitrarily close to the lower bound of the convex function inducing the target function, which is also a lower bound of $\mathfrak{L}$ but does not depend on any concrete neural network. Bad local valleys are harmful to the optimization problem because if the training process enters them when optimizing the target function, the parameters obtained at the end of the learning process are provably not optimal. Nguyen proved that, for activation functions as before that also satisfy that there are no nonzero coefficients $(\lambda_{i},a_{i})_{i=1}^{p}$ with $a_{i}\neq a_{j}$ for all $i\neq j$ such that $\varphi(x)=\sum_{i=1}^{p}\lambda_{i}\varphi(x-a_{i})$ for all $x\in\mathbb{R}$, FCFNNs with a layer $l\in[L-1]$ satisfying $N_{l}>\left|\mathcal{D}_{\text{train}}\right|$ and $N_{i}>N_{i+1}$ for $i\in\{l+1,\ldots,L\}$ induce target functions $\mathfrak{L}$ with no bad local valleys and, if $l\leq L-2$, then every local valley of $\mathfrak{L}$ is unbounded. This implies that from any initial parameter, there is a continuous path on which the loss function is nonincreasing from it to a point that is arbitrarily close to the infimum of the loss.
@@
$\sup_{\theta\in\Theta}\left|\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\phi_{\mathcal{N}_{\theta}})-\mathcal{R}(\phi_{\mathcal{N}_{\theta}})\right|\leq 2B\sqrt{\frac{(\text{dim}_{\text{PH}}\Theta+1)\log^{2}(mL^{2})}{m}+\frac{\log(7M/\gamma)}{m}}$
@@ 4 Challenges, future directions, and conclusions
In this survey we have seen many examples of how topological data analysis, and particularly (persistent) homology and Mapper, can be applied to study the properties of neural networks. In Section 3.1, we saw how to compute two homology groups that yielded different information about a neural network only by taking its graph without weights. However, we noticed that the ranks of those homology groups were invariant to many important structural properties of the neural network, such as the order of the layers. Also, their values were simply a combination of the width and the depth of the graph, not capturing much information about the network. To obtain better insight into the properties of neural networks, more fine-grained homology theories for directed graphs are needed. A promising line of work would be to study the persistent path homology (Chowdhury and Mémoli, 2018) of graphs augmented with the weights associated to each edge, as in Section 3.3.
In Section 3.2, we saw how topology was useful in recovering the topology of decision regions and boundaries, with applications especially in model selection. Furthermore, we saw that GTDA, an evolution of the Mapper algorithm for graphs as input, was successful in analyzing patterns in output spaces of neural networks, leading to a better understanding of misclassification errors in several datasets.
@@
Thanks to the differentiability theory of persistence diagrams, we saw how to improve decision regions by simplifying them. In addition, we discussed how the topology was useful for capturing the quality —measured in different ways— of generative models. Particularly interesting is the study of the disentanglement of generative models, since being able to decouple the different sources of variation could lead to a better control and quality of the networks’ outputs. We find that refining the article by Zhou et al. (2021) and using its measures to regularize generative neural networks could be an interesting approach to improving neural network design using TDA.
@@
In Section 3.3 we analyzed the methods studying the neural network parameters and activations. We split the section into the articles using Mapper and (persistent) homology, because of the differences in their approaches. In the first case, we found both studies of the parameters and of the activations. In the first case, we saw how Carlsson and Brüel Gabrielsson found very interesting connections between the topologies of the space of natural images and the space of weights of some layers in convolutional networks, among many other interesting topology configurations of the convolutional weights. Also, we saw how Gabella studied the Mapper graphs of the evolution of the weights during training, finding a very interesting pattern in the weight distribution of one of the architectures tried in his experiments: the weights seemed to be distributed in a surface! A closer look at the topology of the weight evolution for a higher variety of architectures must be taken to realize if the weights share common patterns in similar problems. This could potentially enable leveraging this structure, as proposed in Carlsson and Brüel Gabrielsson (2020), for instance. For activations, we found that the general approach proposed by TopoAct (Rathore et al., 2021) was extremely useful for understanding the internal behavior of neural networks, with the many ramifications presented in the survey.
@@
Computational problems motivate the development of more suitable algorithms to compute persistence diagrams or alternative invariants that also capture information of neural networks. Although powerful tools for computing Vietoris–Rips persistence diagrams are available, such as Ripser (Bauer, 2021), Ripser++ (Zhang et al., 2020), Gudhi (The GUDHI Project, 2015), SLINK (Sibson, 1973), algorithms that parallelize computations are of special interest due to the increasing availability of powerful hardware and software methods to perform distributed computing. Also, approximating known filtrations by means of new filtrations, as in (Sheehy, 2013), or by means of machine learning models, as in Montufar et al. (2020).
As we have seen in this survey, there are many different ways to produce meaningful filtrations for neural networks in all the domains we reviewed. For this reason, it would be desirable to capture the joint information of some of them at the same time. Also, for some specific problems, like analyzing the evolution of the data distribution through the layers, filtrations by distances alone may not be good enough, as they do not consider important data properties such as the density of points in the sample taken. Luckily, multiparameter persistence (Botnan and Lesnick, 2023) deals with this kind of problem setting, allowing to extract information from non-linear filtrations, that is, with more than one varying parameter. Unluckily, multiparameter persistence does not have a canonical, easy-to-use and computationally efficient representation like persistence diagrams for one-dimensional persistence. However, there are already works proposing a rich body of useful, and even differentiable, representations for multiparameter persistence; see, for example Loiseaux et al. (2023a, b); Xin et al. (2023). We see multiparameter persistence as one of the main lines of work not only for the analysis of deep neural networks, but also for the whole machine learning community, as multiparameter persistence provides sharper ways to extract information from data. However, we are far from having usable representations for real neural network use cases due to the huge quantity of points in real datasets and neurons in architectures. For this reason, further fundamental research on the topic is needed before we can grasp its advantages in the deep learning community. Further work in this direction could be based on three basic pillars: 1. Ease of use of multiparameter representations that can be stored in any conventional computer, since representations must be intuitive for the average machine learning researcher; 2. Efficiency of computation; 3. Differentiability of representations with respect to the point clouds, to allow regularization in learning tasks.
In conclusion, we have seen that the use of TDA applied to studying neural networks is an exciting path that can be useful in many scenarios. TDA, although computationally expensive, is a tool that has been shown to be connected with many interesting properties of neural networks such as generalization in classification problems or entanglement and latent space quality in generative models. Moreover, we have seen that there are many ways in which TDA can be used in applied scenarios, converting TDA into an essential tool for deep learning practitioners.
@@
Appendix A Peer-reviewed articles discussed in this survey
Table 1: Table containing the published articles reviewed in this survey ordered by year of publication. Each row corresponds to an article. The first column contains the title; the second column contains a one-line summary; the third column contains the categories introduced in Section 1 associated to each article; and the last column contains the applications introduced in Section 2.4. The four possible categories are 1. Structure of the neural network; 2. Input and output spaces; 3. Internal representations and activations; 4. Training dynamics and loss functions. The seven possible applications are 1. Regularization; 2. Pruning of neural networks; 3. Detection of adversarial, out-of-distribution and shifted examples; 4. Detection of trojaned networks; 5. Model selection; 6. Prediction of accuracy; 7. Quality assessment of generative models.
Title Summary Cat. Apps.
On the complexity of neural network classifiers: A comparison between shallow and deep architectures (Bianchini and Scarselli, 2014)
Bounds on the Betti numbers of the positive decision region generated by binary classification neural networks with Pfaffian activations
(2) -
Topological approaches to deep learning (Carlsson and Brüel Gabrielsson, 2020)
Topological analysis of the weights of convolutional neural networks and generalization of convolutional neural networks
(3) -
Topological data analysis of decision boundaries with application to model selection (Ramamurthy et al., 2019)
Study and approximation of the topology of network decision boundaries
(2) (5)
A topological regularizer for classifiers via persistent homology (Chen et al., 2019)
Regularization of neural networks by modifying their decision regions using differentiable persistent homology
(2) (1)
Geometry Score: A method for comparing generative adversarial networks (Khrulkov and Oseledets, 2018)
Measurement of GAN quality using persistent homology
(2) (7)
What does it mean to learn in deep networks? And, how does one detect adversarial attacks? (Corneanu et al., 2019)
Study of generalization in terms of the persistent homology of neuron activations and their correlations
(3) (1, 3)
On connected sublevel sets in deep learning (Nguyen, 2019)
Study of the connectivity, boundedness, and local minima of sublevel sets of convex losses for overparameterised neural networks with piecewise linear activation functions
(4) -
Neural persistence: A complexity measure for deep neural networks using algebraic topology (Rieck et al., 2019)
Topological complexity measure for neural networks based on their weights
(3) (1)
Characterizing the shape of activation space in deep neural networks (Gebhart et al., 2019)
Topological characterization of the neuron activations given a sample
(3) (3)
Exposition and interpretation of the topology of neural networks (Brüel Gabrielsson and Carlsson, 2019)
Topological analysis of the weights of convolutional neural networks and connection to generalization capacity of models
(3) -
Path homologies of deep feedforward networks (Chowdhury et al., 2019)
Analysis of path and directed flag homology groups of MLPs’ directed graph
(1) -
Topology of deep neural networks (Naitzat et al., 2020)
Study of the topology of the data through layer transformations
(3) -
Finding the homology of decision boundaries with active learning (Li et al., 2020)
Use of active learning to improve the methods of Topological data analysis of decision boundaries with application to model selection
(2) (5)
Computing the testing error without a testing set (Corneanu et al., 2020)
Regression of test accuracy using topological vectorizations of persistence diagrams
(3) (6)
Topological detection of trojaned neural networks (Zheng et al., 2021)
Study of trojaned networks in terms of the persistent homology of neuron activations and their correlations
(3) (4)
Experimental stability analysis of neural networks in classification problems with confidence sets for persistence diagrams (Akai et al., 2021)
Study of persistence diagrams of last hidden layer activations and its connection with generalization
(3) -
PHom-GeM: Persistent homology for generative models (Charlier et al., 2019)
Comparison between persistence diagrams of real and generated manifolds using generative models
(2) (7)
TopoAct: Visually exploring the shape of activations in deep learning (Rathore et al., 2021)
Analysis of the Mapper graph of neuron activations for each layer
(3) -
Deep neural network pruning using persistent homology (Watanabe and Yamana, 2020)
Pruning of neural networks using persistent homology
(3) (2)
Intrinsic dimension, persistent homology and generalization in neural networks (Birdal et al., 2021)
Generalization error bounds using persistent homology dimension on the training weights
(3, 4) (1)
Activation landscapes as a topological summary of neural network performance (Wheeler et al., 2021)
Study of the topology of the data through layer transformations and its connection with training accuracy
(3) -
Topology of learning in feedforward neural networks (Gabella, 2021)
Analysis of the evolution of the weights of a neural network during training using Mapper
(3) -
Topological uncertainty: Monitoring trained neural networks through persistence of activation graphs (Lacombe et al., 2021)
Uncertainty measurement of neural network predictions using the topology of neuron activations
(3) (3, 5)
Topological measurement of deep neural networks using persistent homology (Watanabe and Yamana, 2022b)
Computation of persistence diagrams using neuron path relevance and connection with network expressivity and problem difficulty
(3) -
Evaluating the disentanglement of deep generative models with manifold topology (Zhou et al., 2021)
Measurement of disentanglement of generative neural networks
(2) -
Quantitative performance assessment of CNN units via topological entropy calculation (Zhao and Zhang, 2022)
Measurement of quality of convolutions in a neural network using persistent homology
(3) -
Overfitting measurement of deep neural networks using no data (Watanabe and Yamana, 2022a)
Study of overfitting by analysing the points near the diagonal of a persistence diagram generated from the weights of a neural network
(3) -
An adversarial robustness perspective on the topology of neural networks (Goibert et al., 2022)
Analysis of adversarial examples by means of the topology of the subgraph of under-optimized edges of neural networks
(3) (3)
Representation topology divergence: A method for comparing neural network representations (Barannikov et al., 2022)
Definition of similarity between data representations using persistent homology
(3) (7)
On the topological expressive power of neural networks (Petri and Leitão, 2020)
Measurement of expressivity of network architectures
(2) -
Generalization bounds using data-dependent fractal dimensions (Dupuis et al., 2023)
Data-dependent generalization error bounds using persistent homology dimension on the training weights
(3, 4) -
TopoBERT: Exploring the topology of fine-tuned word representations (Rathore et al., 2023)
Mapper graph of transformer-based models with applications in finetuning of language models
(3) -
Experimental observations of the topology of convolutional neural network activations (Purvine et al., 2023)
Analysis of the Mapper graph of neuron activations and definition of a similarity function between layers using sliced Wasserstein distances between persistence diagrams generated from them
(3) -
ReLU neural networks, polyhedral decompositions, and persistent homology (Liu et al., 2023b)
Detection of homological features in manifolds embedded in the input space of a neural network using the polyhedra decomposition induced by a ReLU feedforward neural network
(2) -
Caveats of neural persistence in deep neural networks (Girrbach et al., 2023)
Connection between neural persistence and variance measures of neural network weights
(3) -
On the use of persistent homology to control the generalization capacity of a neural network (Barbara et al., 2024)
Regression of generalization gap using the average persistence of zero dimensional persistence diagrams
(3) (6)
Visualizing and analyzing the topology of neuron activations in deep adversarial training (Zhou et al., 2023)
Analysis of the Mapper graph of neuron activations of normally and adversarially trained neural networks
(3) -
TopP&R: Robust Support Estimation Approach for Evaluating Fidelity and Diversity in Generative Models (Kim et al., 2023)
Approximation of the precision and recall scores for generative models using a persistent homology-based approach.
(2) (7)
Topological structure of complex predictions (Liu et al., 2023a)
Analysis of the GTDA Reeb network of output spaces of neural networks
(2) -
Topological dynamics of functional neural network graphs during reinforcement learning (Muller et al., 2024)
Study of reinforcement learning neural networks during training and inference using homology
(3) -
Appendix B The Mapper algorithm
Algorithm 1 Mapper algorithm
0:  Simplicial complex $S_{\mathcal{D}}$.
1:  $S_{\mathcal{D}}\leftarrow\emptyset$
3:  for all $i\in I$ do
6:  end for
8:     if $\bigcap_{j=1}^{t}C_{j}\neq\emptyset$ then
10:     end if
11:  end for
12:  return  $S_{\mathcal{D}}$